Your project will give you a rate of return of 100% over 15 years. Is this a lot or a little? How does this 15-year rate of capital accumulation compare to a rate of 1% over 3 months?
Answer #1

Rate = 100% for 15 years
for Annual rate

(1+r)15 -1 = 100%
r = 21/15-1 = 0.4729 or 4.73% It is little as compared to stock returns and mutual funds returns.

Since rate = 1% for 3 months

Number of periods = 15*12/3 = 60

The rate for 15 years id rate of 1% for 3 months = (1+r)60-1 = (1+1%)60-1 = 81.67%

Hence the 15 year accumulation is more as compared to 1% rate over 3 months.

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well